# Gravity from entropy

### v4

@@ III.3 Introduction of G-field $\tilde{\mathcal{G}}$ and the field $\tilde{\Theta}$
In order to investigate the property of the modified gravity emerging from the entropic action, here we introduce the auxiliary G-field $\tilde{\bm{\mathcal{G}}}$ and the auxiliary field $\tilde{\bm{\Theta}}$. In this way we will be able to show in the following that the proposed entropic action can be interpreted as the sum of a dressed Einstein-Hilbert action with an emergent positive cosmological constant only depending on the G-field $\tilde{\bm{\mathcal{G}}}$ and a matter action.
@@
This Lagrangian is now linear in $\tilde{\bf G}\tilde{g}^{-1}$, thus by introducing the auxiliary fields and giving a physical meaning to them as encoding for new particles, we have reduced our theory to a theory linear in $\tilde{\bm{\mathcal{R}}}$. The fields $\tilde{\bm{\Theta}}$ and $\tilde{\bm{\mathcal{G}}}$ are given by
@@ III.5.1 Variation with respect to $\tilde{\bm{\mathcal{G}}}$
Note that these are equations determine the relation between the fields $\tilde{\bm{\Theta}}$ the curvature $\tilde{\bm{\mathcal{R}}}$ and the matter field determining $\tilde{\bf M}$. We can write these equations by separating the contribution corresponding to the metric for 00-forms, $1$-forms and $2$-forms obtaining
@@ III.5.3 Variation with respect to ${g}$
${R}^{\mathcal{G}}_{(\mu\nu)}-\frac{1}{2}{g}_{\mu\nu}\Big{(}\mathcal{R}_{\mathcal{G}}-2\Lambda_{\mathcal{G}})+{{\mathcal{D}}}_{(\mu\nu)}=\kappa{{\mathcal{T}}_{(\mu\nu)}},$ (67)
@@ IV Conclusions
This work proposes a modified theory of gravity emerging from statistical mechanics and information theory action. The fundamental idea of this theory is to associate the metric to a density matrix. In particular two metrics are discussed: the metric of spacetime that fully defines its geometry and the metric induced by matter-fields that are effectively curving the space. The interplay between geometry of spacetime and matter fields is explicitly captured by the proposed action given by the quantum relative entropy between the metric and the metric induced by the matter fields.
@@
Here a Lorentzian theory consistent with this fundamental physical interpretation of gravity is formulated. In order to do this we have built the necessary mathematical background to define the entropy and the cross-entropy of metric tensors between two $n$-forms in a Lorentzian well defined way. A modified gravity is emerging from this framework when the matter fields are described by topological Dirac-Kähler bosons formed by the direct sum between a 0-form, a 1-form and a 2-form and the induced metric also depends on the curvature of spacetime. The modified Einstein equations reduce to the Einstein equations in the regime of low coupling. By introducing the G-field we obtain the modified Einstein equations and the equations of motion for the matter and the G-field. From this theory it emerges that the proposed entropic action takes the form of a sum between a dressed Einstein-Hilbert action and a matter action. Interesting the dressed Einstein-Hilbert action displays an emergent positive cosmological constant that only depends on the G-field. Moreover, thanks to the introduction of the G-field, the equations of modified gravity remains second order in the metric, in the matter and in the G-field.
The interpretation of the metric as a density matrix is shown to be very useful, albeit we have deformed a bit the notion of density matrices relaxing the constraints of having unitary trace, and requiring the existence of the inverse. These choices are motivated here by the necessity to have a Lorentzian invariant theory. Moreover this generalization of the notion of density matrices points out to possible connections of this approach to von Neumann algebras and the theory of entanglement in quantum field theories [17] that will be explored in future works.
@@
In conclusion, we hope that this approach can help identify the deep connections between gravity and statistical physics. Given the interpretation of the metric as a density matrix our hope is that this approach will also be instrumental to formulate new approaches to quantum gravity. Finally future investigation might explore the role of the the G-field in dark matter. Future directions in this line of research involve also the investigation of the proposed entropic action under the renormalization group, and possible connections with phenomenology and experimental results.
@@
Appendix A Quantum interpretation: q-forms
In order to outline a possible quantum interpretation of our theory we consider here a toy quantum example that extends the notion of a q-bits to q-forms. We assume that q-forms belong to the Hilbert space $\mathcal{H}_{p}^{(1)}$ defined at point $p$ of coordinates ${\bf x}$ of the considered $d$-dimensional manifold $\mathcal{K}$. This Hilbert space includes as states all q-forms or more specifically “quantum 1-forms” |$\ket{\omega_{p}}$ defined at $p$. A given choice of coordinate systems for our manifold define a basis for the Hilbert space $\mathcal{H}_{p}^{(1)}$ with basis |${\ket{dx^{\mu}}}$ with $\mu\in\{1,2,\ldots,d\}$.
Thus quantum $1$-form |$\ket{\omega}$ is given by
|$\ket{\omega_{p}}=\omega_{\mu}({\bf x})\ket{dx^{\mu}},$ (70)
where $\omega_{\mu}({\bf x})$ is the probability amplitude that the q-form is along $dx^{\mu}$. A possible quantum interpretation of the metrics for $1$-forms $g$ and ${\bf G}$, is then as a “renormalizable density matrices” for q-forms.
This construction of q-forms can be similarly extended to higher-order q-forms such as q-forms of order 2. These latter q-forms live in the Hilbert space $\mathcal{H}_{p}^{(2)}$ of with basis |$\ket{dx^{\mu}\wedge dx^{\mu}}$ with $\mu,\nu\in\{1,2,\ldots,d\}$ and $\mu<\nu$ with
[…]
Consequently the $q$-form |$\ket{\zeta_{p}}$ can be expressed as
|$\ket{\zeta_{p}}=\zeta_{\mu\nu}({\bf x})\ket{dx^{\mu}\wedge dx^{\mu}}.$ (72)
Accordingly, a possible quantum interpretation of the metric for $2$-forms $g_{(2)}$ or ${\bf G}_{(2)}$, defined in Sec. III is then as “renormalizable density matrices” for q-forms of order $2$.
It is to note that a full topological theory of q-forms involving at the same time 00, $1$ and $2$ q-forms will need to be compatible with the uncertainty principle. Thus this theory might involve non trivial fundamental assumptions on the quantum topological q-forms which are beyond the scope of this work.
Appendix B The effect of renormalizing the metric matrices pointwise
Let assume for the moment that instead of considering the metric matrices $g$ and ${\bf G}$, as “renormalizable density matrices” at point $p$, we normalize $g$ and ${\bf G}$ at each point in space. In this scenario we would consider the density matrices $\rho_{g}$ and $\sigma_{\bf G}$ given by
$\rho_{g}=\frac{g}{\mbox{Tr}g},\quad\sigma_{\bf G}=\frac{\bf G}{\mbox{Tr}{\bf G}}.$ (73)
The first observation that we make is that with this choice, we cannot any longer treat $\rho_{g}$ and $\rho_{g}^{-1}$ on the same footing, as we do for $g$ and $g^{-1}$ as $\rho_{g}$ will have unitary trace while $\rho_{g}^{-1}$ will not.
Another difference emerges if we want to consider the dynamics of the warm-up scenario in which we use $\rho_{g}$ and $\sigma_{\bf G}$ instead of directly $g$ and ${\bf G}$. Indeed the Lagrangian $\mathcal{L}$ defined in Eq.(10) would turn into $\hat{\mathcal{L}}$
^$\hat{\mathcal{L}}=\mbox{Tr}\rho_{g}\ln{\rho_{g}^{-1}}-\mbox{Tr}\rho_{g}\ln\sigma_{G}^{-1},$ (74)
where the Lorentz invariant trace for rank-$2$ tensor is defined as in Sec. II.1. Let us now investigate in the warm-up example, the consequence of renormalizing the metric matrices in this way. We observe that since $g$ has all eigenvalue equal to one, its trace is simply $\mbox{Tr}g=d$, thus
$\rho_{g}=\frac{g}{d},$ (75)
and the entropy $\hat{H}$ becomes non zero and given by $\ln d$. In fact, using $H=\mbox{Tr}g\ln g^{-1}=0$ we have
$\hat{H}$ $=$ $\mbox{Tr}\rho_{g}\ln(\rho_{g})^{-1}=\ln d.$ (76)
If we take ${\bf G}$, the metric induced by the matter field as in Eq.(14), we have than that its trace is given by
$\mbox{Tr}{\bf G}=d+\alpha|\nabla\phi|^{2}.$ (77)
Using these results it can be easily shown that we can express the Lagrangian $\hat{\mathcal{L}}$ using the normalized density matrices $\rho_{g}$ and $\sigma_{G}$ and given by Eq.(74) as a function of $\mathcal{L}$ given by Eq. (10), i.e.
^$\hat{\mathcal{L}}=\frac{1}{d}\mathcal{L}+\ln\Big{(}1+\frac{\alpha}{d}|\nabla\phi|^{2}\Big{)}.$ (78)
Using the Eq.(20) to express $\mathcal{L}$ we thus obtain
^$\hat{\mathcal{L}}=-\frac{1}{d}\ln\Big{(}1+\alpha|\nabla\phi|^{2}\Big{)}+\ln\Big{(}1+\frac{\alpha}{d}|\nabla\phi|^{2}\Big{)}.$ (79)
If follow that for $\alpha^{\prime}\ll 1$ we obtain that the first non-trivial term is given by
^$\hat{\mathcal{L}}=\frac{d-1}{2d^{2}}\alpha^{2}(|\nabla\phi|^{2})^{2}.$ (80)
In other words, in this limit we will not recover the zero-mass Klein-Gordon action as we obtain with the action $\mathcal{L}$ using directly $g$ and ${\bf G}$ instead of $\rho_{g}$ and $\sigma_{\bf G}$. Therefore this exercise provides us an hint of the relevance to use the metric matrices as “renormalizable density matrices”.
Appendix C Eigenvalues and entropy for metrics between 1-forms and for metric between $n$-forms
@@
In this Appendix our goal is to use algebraic geometry [49] to define the eigenvalues of the metrics between $n$-forms and their associated entropy. In order to establish the foundation of this treatment let us revisit the treatment outlined in Sec. II for defining the eigenvectors and the associated entropy of the metric for $1$-forms and treat on the same footing metrics tensor between pair of $n$-forms with arbitrary $n$.
C.0.1 Eigenvalues of metric matrices ${\bf G}_{(n)}$
@@
$\omega=\omega_{\rho}dx^{\rho}$ (81)
and the metric for among two of such $1$-forms can be written as
@@
where ${[G_{(1)}]}_{\mu\nu}$ is Hermitian. The contraction between the metric ${\bf G}_{(1)}$ and the generic one form $\omega$ is defined as
@@
The eigenvalue problem of a metric ${\bf G}_{(1)}$ among $1$-forms is then defined as
@@
We are now in the position to extent the formalism to the metricx ${\bf G}_{(2)}$ among two forms of order two. The generic two form $\zeta$ is defined as
@@
The metric ${\bf G}_{(2)}$ for two-forms is defined as
@@
In total analogy with the previous case, the eigenvalue problem associated to the metric ${\bf G}_{(2)}$ among $2$-forms, is defined as
@@
Let us now consider the particular case in which ${\bf G}_{(2)}={g}_{(2)}$. In this case the eigenvalue problem Eq.(98) reads
@@
The treatment of the eigenvalues associated to the metrics tensors ${\bf G}_{(n)}$ between $n$-forms can be performed in a straightforward way, following similar steps. Let us observe that the generic $n$-form is given by
@@
where ${\bf G}_{(n)}$ is antisymmetric in the first $n$ indices and in the last $2$ indices and Hermitian under the exchange of the first $n$ indices with the second $n$ indices.
@@
Here and in the following we will use the notation ${[g_{(n)}]}_{\vec{\mu}\vec{\nu}}$ of indicating the element of this tensor where $\vec{\mu}=(\mu_{1}\mu_{2}\ldots\mu_{n})$ and $\vec{\nu}=(\nu_{1}\nu_{2}\ldots\nu_{n})$ indicate $n$-tuples of indices. Using an analogous notation, we define the scalar product between the generic metric ${\bf G}_{(n)}$ between pairs of $n$-forms and the $n$ form $\zeta$ as as
@@
C.0.2 Flattening of the metric tensors ${\bf G}_{(n)}$ and reduction to a matrix eigenvalue problem
@@
The eigenvalue problem for 2-forms Eq.(98) can be as well written in matrix form, by considering the $m\times m$ flattened matrices ${\bf G}_{(2)}^{F}$ and $g_{(2)}^{F}$ with $m=d(d-1)/2$ representing the dimension of the basis for the 2-forms defined on the $d$ dimensional manifold $\mathcal{K}$. For instance, taking the basis $\{dx^{\mu}\wedge dx^{\nu}\}$ with $\mu<\nu$ the flattened matrices ${\bf G}_{(2)}^{F}$ and $g_{(2)}^{M}$ will have respectively matrix elements
@@
Similarly metric tensors ${\bf G}_{(n)}$ between $n$-forms can be flattened by considering a basis of independent $n$-forms. For instance a basis for these $n$ forms can be given by the canonical $n$ forms
@@
C.0.3 Trace of the metric tensors ${\bf G}_{(n)}$

### v5

@@ III.3 Introduction of G-field $\tilde{\mathcal{G}}$ and the field $\tilde{\Theta}$
In this section our goal is to investigate the property of the modified gravity emerging from the entropic action. To this end we introduce the auxiliary G-field $\tilde{\bm{\mathcal{G}}}$ and the auxiliary field $\tilde{\bm{\Theta}}$. In this way we will be able to show in the following that the proposed entropic action can be interpreted as the sum of a dressed Einstein-Hilbert action with an emergent positive cosmological constant only depending on the G-field $\tilde{\bm{\mathcal{G}}}$ and a matter action.
@@
This Lagrangian is now linear in $\tilde{\bf G}\tilde{g}^{-1}$. Note that a conservative interpretation of this transformation will not give a physical interpretation to the auxiliary fields. Thus, according to this point of view, the auxiliary fields are not changing the physics of the original Lagrangian and this theory might be prone to Ostrogradsky instability. However, we known from statistical physics that Lagrangian multipliers can also acquire a rather physical meaning. For instance the temperature and the chemical potential are Lagrangian multipliers that are measurable. Here we would like to embrace this point of view and give a physical meaning to the introduced Lagrangian multipliers as physical fields defined on our manifold $\mathcal{K}$. Interpreting in this way the Lagrangian multipliers changes the physics of the problem as the phase space associated to the equations of motion acquire new dimensions and the fixed points as well as the initial value problem will involve the newly introduced fields. In this latter interpretation, introducing the auxiliary fields and giving a physical meaning to them as encoding for new fields, we have reduced our theory to a theory linear in $\tilde{\bm{\mathcal{R}}}$. In particular, as we will see below, the equations of motion will involve at most second order derivatives of the fields, thus the model will avoid the Ostrogradsky instability.
The fields $\tilde{\bm{\Theta}}$ and $\tilde{\bm{\mathcal{G}}}$ are given by
@@ III.5.1 Variation with respect to $\tilde{\bm{\mathcal{G}}}$
Note that these are the equations that determine the relation between the fields $\tilde{\bm{\Theta}}$ the curvature $\tilde{\bm{\mathcal{R}}}$ and the matter field determining $\tilde{\bf M}$. We can write these equations by separating the contribution corresponding to the metric for 00-forms, $1$-forms and $2$-forms obtaining
@@ III.5.3 Variation with respect to ${g}$
${R}^{\mathcal{G}}_{(\mu\nu)}-\frac{1}{2}{g}_{\mu\nu}\Big{(}\mathcal{R}_{\mathcal{G}}-2\Lambda_{\mathcal{G}}\Big{)}+{{\mathcal{D}}}_{(\mu\nu)}=\kappa{{\mathcal{T}}_{(\mu\nu)}},$ (67)
@@ IV Conclusions
This work proposes a modified theory of gravity emerging from statistical mechanics and information theory action. The fundamental idea of this theory is to associate the metric to a quantum operator, playing the role of a renormalizable and effective density matrix. In particular two metrics are discussed: the metric of spacetime that fully defines its geometry and the metric induced by matter-fields that are effectively curving the space. The interplay between geometry of spacetime and matter fields is explicitly captured by the proposed action given by the quantum relative entropy between the metric and the metric induced by the matter fields.
@@
Here a Lorentzian theory consistent with this fundamental physical interpretation of gravity is formulated. In order to do this we have built the necessary mathematical background to define the entropy and the cross-entropy of metric tensors in a Lorentzian well defined way. A modified gravity is emerging from this framework when the matter fields are described by topological Dirac-Kähler bosons formed by the direct sum between a 0-form, a 1-form and a 2-form and the induced metric also depends on the curvature of spacetime. The modified Einstein equations reduce to the Einstein equations in the regime of low coupling. By introducing the G-field we obtain the modified Einstein equations and the equations of motion for the matter and the G-field. From this theory it emerges that the proposed entropic action takes the form of a sum between a dressed Einstein-Hilbert action and a matter action. Interestingly, the dressed Einstein-Hilbert action displays an emergent positive cosmological constant that only depends on the G-field. Moreover, thanks to the introduction of the G-field, the equations of modified gravity remains second order in the metric, in the matter and in the G-field.
The interpretation of the topological metrics tensors as a quantum operators, or effective density matrix where we have relaxed the constraints of having unitary trace, and we have required the existence of the inverse is shown to be very useful. These choices are motivated here by the necessity to have a Lorentzian invariant theory. Here we have established the relation of the adopted quantum relative action with Araki quantum relative entropy [37], opening the perspective of using the theory of quantum von Neumann algebras and the theory of entanglement [17] to investigate further the properties of the proposed theory.
@@
In conclusion, we hope that this approach can help identify the deep connections between gravity, quantum mechanics and statistical physics. Given the interpretation of the metrics as a quantum operators our hope is that this approach will also be instrumental to formulate approaches to quantum gravity in second quantization. Finally future investigation might explore the role of the the G-field in dark matter. Future directions in this line of research involve also the investigation of the proposed entropic action under the renormalization group, and possible connections with phenomenology and experimental results.
@@
Appendix A Inclusion of an Abelian gauge field
An Abelian gauge field $A_{\mu}$ can be easily included in the general framework introduced in this work. We consider the operator $\tilde{\bf F}$ defined as
$\tilde{\bf F}=0\oplus 0^{\mu}dx_{\mu}\oplus F_{\mu\nu}F_{\rho\sigma}\Big{(}dx^{\mu}\wedge dx^{\nu}\Big{)}\otimes\Big{(}dx^{\rho}\wedge dx^{\sigma}\Big{)},$
where $F_{\mu\nu}=\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$ and we consider the following expression for the metric induced by matter fields
$\tilde{\bf G}$ $=$ $\tilde{g}+\alpha\Big{(}{\tilde{\bf M}+\tilde{\bf F}}\Big{)}-\beta{\tilde{\bm{\mathcal{R}}}}.$ (70)
Note that in this expression we modify as well the definition of $\tilde{\bf M}$ as the Dirac operator will be defined as in Eq.(32) by performing the minimal substitution in the covariant derivative
$\nabla_{\mu}\to\nabla_{\mu}^{(A)}=\nabla_{\mu}-\textrm{i}eA_{\mu}.$ (71)
Appendix B Eigenvalues and entropy for metrics ${\bf G}_{(1)}$ and for metrics ${\bf G}_{(n)}$
@@
In this Appendix our goal is to use algebraic geometry [54] to define the eigenvalues of the metrics ${\bf G}_{(n)}$ considered as quantum opeartors between $n$-forms and their associated entropy. In order to establish the foundation of this treatment let us revisit the treatment outlined in Sec. II for defining the eigenvectors and the associated entropy of the metrics ${\bf G}_{(1)}$ viewed as qauntum operators between $1$-forms and treat on the same footing metrics tensor ${\bf G}_{(n)}$ viewed as quantum operators between $n$-forms with arbitrary values of $n$.
B.1 Eigenvalues of metric matrices ${\bf G}_{(n)}$
@@
$\omega=\omega_{\rho}dx^{\rho}.$ (72)
The metric ${\bf G}_{(1)}$ can be interpreted as a quantum operator among two of such $1$-forms and can be written as
@@
where ${[G_{(1)}]}_{\mu\nu}$ is Hermitian. The $\cdot$ product between the metric ${\bf G}_{(1)}$ and the generic one form $\omega$ is defined as
@@
The eigenvalue problem of a metric ${\bf G}_{(1)}$ is then defined as
@@
Note that ${[{\bf G}_{(1)}]}_{\mu\nu}$ admits also an interpretation as metric between vectors. We are now in the position to extent the formalism to the metrics ${\bf G}_{(2)}$ considered as quantum operators among two forms of order two. The generic two form $\zeta$ is defined as
@@
The metric ${\bf G}_{(2)}$ considered as a quantum operator between two two-forms is defined as
@@
In total analogy with the previous case, the eigenvalue problem associated to the metric ${\bf G}_{(2)}$ viewed as quantum operator among two $2$-forms, is defined as
@@
Note that ${[{\bf G}_{(2)}]}_{\mu\nu\rho\sigma}$ admits also an interpretation as metric between bivectors. Let us now consider the particular case in which ${\bf G}_{(2)}={g}_{(2)}$. In this case the eigenvalue problem Eq.(89) reads
@@
The treatment of the eigenvalues associated to the metrics tensors ${\bf G}_{(n)}$ considered as quantum operators between two $n$-forms can be performed in a straightforward way, following similar steps. Let us observe that the generic $n$-form is given by
@@
where ${\bf G}_{(n)}$ is antisymmetric in the first $n$ indices and in the last $n$ indices and Hermitian under the exchange of the first $n$ indices with the second $n$ indices.
@@
Here and in the following we will use the notation ${[g_{(n)}]}_{\vec{\mu}\vec{\nu}}$ of indicating the element of this tensor where $\vec{\mu}=(\mu_{1}\mu_{2}\ldots\mu_{n})$ and $\vec{\nu}=(\nu_{1}\nu_{2}\ldots\nu_{n})$ indicate $n$-tuples of indices. Using an analogous notation, we define the $\cdot$ product between the generic metric ${\bf G}_{(n)}$ and the $n$-form $\zeta$ as
@@
B.2 Flattening of the metric tensors ${\bf G}_{(n)}$ and reduction to a matrix eigenvalue problem
@@
The eigenvalue problem for the metric ${\bf G}_{(2)}$ Eq.(89) can be as well written in matrix form, by considering the $m\times m$ flattened matrices ${\bf G}_{(2)}^{F}$ and $g_{(2)}^{F}$ with $m=d(d-1)/2$ representing the dimension of the basis for the 2-forms defined on the $d$ dimensional manifold $\mathcal{K}$. For instance, taking the basis $\{dx^{\mu}\wedge dx^{\nu}\}$ with $\mu<\nu$ the flattened matrices ${\bf G}_{(2)}^{F}$ and $g_{(2)}^{M}$ will have respectively matrix elements
@@
Similarly metric tensors ${\bf G}_{(n)}$ considered as quantum operators between $n$-forms can be flattened by considering a basis of independent $n$-forms. For instance a basis for these $n$ forms can be given by the canonical $n$-forms
@@
B.3 Trace of the metric tensors ${\bf G}_{(n)}$